# RACE-IT: A Reconfigurable Analog Computing Engine for In-Memory Transformer Acceleration

### v2

@@ Abstract
We compare RACE-IT with various accelerators. Results show that RACE-IT increases performance by 453$\times$ and 15$\times$, and reduces energy by 354$\times$ and 122$\times$ over the state-of-the-art GPUs and existing Transformer-specific IMC accelerators, respectively.
@@ II-A Crossbar-based VMM Computing Engine
RRAM is an emerging non-volatile memory technology with strong potential for accelerating VMMs when arranged in crossbar arrays [19]. Figure 1 illustrates a VMM example where a 3-element input feature vector ($I$) is multiplied by a $\times 3$ 3 weight matrix ($W$) using a RRAM crossbar. Digital-to-Analog Converters (DACs) convert input elements into a voltage vector ($v$), which is applied to the word lines. Weights are stored as RRAM conductance values ($g$), and Kirchhoff’s law ensures that the resulting current ($i$) on each bit line reflects the dot product of the input voltages and the conductance values. These analog currents are sampled by Sample-and-Hold (S&H) units and then digitized by Analog-to-Digital Converters (ADCs) to produce the digital output vector ($O$). While the figure shows a small $\times 3$ 3 crossbar for clarity, practical systems often use multiple much larger arrays (e.g., 512×256 [18]) to exploit massive parallelism.
@@ II-B Analog Content Addressable Memory
Content Addressable Memories (CAMs) are a memory structure ubiquitous in networking, in which an input data vector is searched in the memory and its location is returned as output [20]. The recently proposed ACAM cell structure [10] allows comparing an analog input value within a defined range stored in the cell. Figure 2(a) presents the schematic of an ACAM cell. Two RRAMs, $M_{1}$ and $M_{2}$, are utilized to store the lower and upper bounds of a range, respectively. The blue-shaded component compares the input $X$ on the Data Line ($DL$) with the lower bound $L$ of the range, while the red-shaded component compares the input $X$ with the upper bound $U$ of the range. The combined operation of the series transistors ($S_{1}$ and $S_{2}$) and RRAMs acts as voltage dividers, controlling the voltage supplied to the pull-down transistors ($T_{1}$ and $T_{2}$). In the blue-shaded component, when the input is greater than or equal to the data stored in $M_{1}$, $T_{1}$ is off, thus keeping the Matching Line ($M$L) at a high voltage. The red-shaded component operates similarly, with an inverter ($T_{3}$ and $T_{4}$) added before the pull-down transistor $T_{2}$. Figure 2(b) illustrates the operational principle of an ACAM cell. It is possible to represent a ”Don’t Care” value on either side of the range by programming $M_{1}$ ($M_{2}$) to the lowest (highest) possible conductance, as depicted by the lower two rows in Figure 2(b). Studies have demonstrated up to 4-bit analog search in ACAM [10, 11].
@@ III-A Cell Structure
Figure 4(a) shows the proposed RACE cell structure. Recall that an ACAM cell in Figure 2(a) consists of a blue component that holds the lower bound of the stored range, and a red component that holds the upper bound. A single RACE cell comprises three blue components and three red components. All six components are connected to the pull-down path of the $ML$. The top two blue components are connected in parallel, implementing a logical AND operation: the $ML$ is pulled down to ground when both $(DL_{1}>M_{1})$ $and$ $(DL_{2}>M_{2})$. Similarly, the bottom two red components are also connected in parallel, so the $ML$ is pulled down when both $(DL_{5}<M_{5})$ $and$ $(DL_{6}<M_{6})$. These two AND combinations are then connected in series with the serially connected blue and red components at the center, which implement the conditions $(DL_{3}>M_{3})$ $OR$ $(DL_{4}<M_{4})$. As a result, the overall operation performed by a RACE cell can be expressed as:
@@ IV-A Activation
By focusing on a specific output bit, we can identify the input ranges that result in that bit being set to 1. For example, output bit $y_{1}$ is 1 when $-0.65\leq x<0$ or $x\geq 0.65$. Figure 5(c) illustrates how these input ranges are mapped to a RACE array. The RACE cells are configured in the one-variable 8-bit comparison mode, with each row responsible for computing a specific output bit. Cells marked with a dash are programmed to cover the full range and thus do not affect the result.
@@ VIII-B Design Space Exploration and Trade-offs
We then evaluate the Energy Delay Product (EDP, lower is better) across depths. Although deeper encoding requires more decoding rounds, the overhead is minimal due to the simplicity of the XOR-based decoding logic. Thus, EDP is mainly affected by additional energy caused by a larger array size. As shown by the yellow line in Figure 11(a), the smallest array size (at depth 3) also achieves the lowest EDP, making it the most energy-efficient configuration. Therefore, we use a Gray code depth of 3 and set the size of each RACE array to 8×508$\times$ 508 × 50 in all subsequent experiments.
@@ VIII-C Performance, Energy and Scalability
Figure 11(b) and (c) show the speedup and energy savings of RACE-IT and other accelerators relative to the GPU baseline. For single-sample inference (BS = 1), which is critical for low-latency, energy-efficient edge AI inference, GPU performance suffers from low utilization [26]. In contrast, RACE-IT achieves 453$\times$ speedup and 354$\times$ energy savings on average. Its advantage over other IMC accelerators stems from replacing costly VFUs and natively supporting attention mechanisms.

### v3

@@ Abstract
[…]
@@ II-A Crossbar-based VMM Computing Engine
RRAM is an emerging non-volatile memory technology with strong potential for accelerating VMMs when arranged in crossbar arrays [19]. Figure 1 illustrates a VMM example where a 3-element input feature vector ($I$) is multiplied by a $3\times 3$ weight matrix ($W$) using a RRAM crossbar. Digital-to-Analog Converters (DACs) convert input elements into a voltage vector ($v$), which is applied to the word lines. Weights are stored as RRAM conductance values ($g$), and Kirchhoff’s law ensures that the resulting current ($i$) on each bit line reflects the dot product of the input voltages and the conductance values. These analog currents are sampled by Sample-and-Hold (S&H) units and then digitized by Analog-to-Digital Converters (ADCs) to produce the digital output vector ($O$). While the figure shows a small $3\times 3$ crossbar for clarity, practical systems often use multiple much larger arrays (e.g., 512×256 [18]) to exploit massive parallelism.
@@ II-B Analog Content Addressable Memory
Content Addressable Memories (CAMs) are a memory structure ubiquitous in networking, in which an input data vector is searched in the memory and its location is returned as output [20]. The recently proposed ACAM cell structure [10] allows comparing an analog input value within a defined range stored in the cell. Figure 2(a) presents the schematic of an ACAM cell. Two RRAMs, $M_{1}$ and $M_{2}$, are utilized to store the lower and upper bounds of a range, respectively. The blue-shaded component compares the input $X$ on the Data Line ($DL$) with the lower bound $L$ of the range, while the red-shaded component compares the input $X$ with the upper bound $U$ of the range. The combined operation of the series transistors ($S_{1}$ and $S_{2}$) and RRAMs acts as voltage dividers, controlling the voltage supplied to the pull-down transistors ($T_{1}$ and $T_{2}$). In the blue-shaded component, when the input is greater than or equal to the data stored in $M_{1}$, $T_{1}$ is off, thus keeping the Matching Line (MML) at a high voltage. The red-shaded component operates similarly, with an inverter ($T_{3}$ and $T_{4}$) added before the pull-down transistor $T_{2}$. Figure 2(b) illustrates the operational principle of an ACAM cell. It is possible to represent a ”Don’t Care” value on either side of the range by programming $M_{1}$ ($M_{2}$) to the lowest (highest) possible conductance, as depicted by the lower two rows in Figure 2(b). Studies have demonstrated up to 4-bit analog search in ACAM [10, 11].
@@ III-A Cell Structure
Figure 4(a) shows the proposed RACE cell structure. Recall that an ACAM cell in Figure 2(a) consists of a blue component that holds the lower bound of the stored range, and a red component that holds the upper bound. A single RACE cell comprises three blue components and three red components. All six components are connected to the pull-down path of the $ML$. The top two blue components are connected in parallel, implementing a logical AND operation: the $ML$ is pulled down to ground when both $(DL_{1}>M_{1})$ andand $(DL_{2}>M_{2})$. Similarly, the bottom two red components are also connected in parallel, so the $ML$ is pulled down when both $(DL_{5}<M_{5})$ andand $(DL_{6}<M_{6})$. These two AND combinations are then connected in series with the serially connected blue and red components at the center, which implement the conditions $(DL_{3}>M_{3})$ OROR $(DL_{4}<M_{4})$. As a result, the overall operation performed by a RACE cell can be expressed as:
@@ IV-A Activation
By focusing on a specific output bit, we can identify the input ranges that result in that bit being set to 1. For example, output bit $y_{1}$ is 1 when $0.65\leq x<0$ or $x\geq 0.65$. Figure 5(c) illustrates how these input ranges are mapped to a RACE array. The RACE cells are configured in the one-variable 8-bit comparison mode, with each row responsible for computing a specific output bit. Cells marked with a dash are programmed to cover the full range and thus do not affect the result.
@@ VIII-B Design Space Exploration and Trade-offs
We then evaluate the Energy Delay Product (EDP, lower is better) across depths. Although deeper encoding requires more decoding rounds, the overhead is minimal due to the simplicity of the XOR-based decoding logic. Thus, EDP is mainly affected by additional energy caused by a larger array size. As shown by the yellow line in Figure 11(a), the smallest array size (at depth 3) also achieves the lowest EDP, making it the most energy-efficient configuration. Therefore, we use a Gray code depth of 3 and set the size of each RACE array to $8\times 50$ in all subsequent experiments.
@@ VIII-C Performance, Energy and Scalability
[…]